\begin{thebibliography}{10}

\bibitem{wolpert1999introduction}
David~H Wolpert and Kagan Tumer.
\newblock An introduction to collective intelligence.
\newblock {\em arXiv preprint cs/9908014}, 1999.

\bibitem{levin2023bioelectric}
Michael Levin.
\newblock Bioelectric networks: the cognitive glue enabling evolutionary scaling from physiology to mind.
\newblock {\em Animal Cognition}, 26(6):1865--1891, 2023.

\bibitem{ozmen2023six}
Ozlem Ozmen~Garibay, Brent Winslow, Salvatore Andolina, Margherita Antona, Anja Bodenschatz, Constantinos Coursaris, Gregory Falco, Stephen~M Fiore, Ivan Garibay, Keri Grieman, et~al.
\newblock Six human-centered artificial intelligence grand challenges.
\newblock {\em International Journal of Human--Computer Interaction}, 39(3):391--437, 2023.

\bibitem{levin2019computational}
Michael Levin.
\newblock The computational boundary of a “self”: developmental bioelectricity drives multicellularity and scale-free cognition.
\newblock {\em Frontiers in psychology}, 10:2688, 2019.

\bibitem{levin2022technological}
Michael Levin.
\newblock Technological approach to mind everywhere: an experimentally-grounded framework for understanding diverse bodies and minds.
\newblock {\em Frontiers in systems neuroscience}, 16:768201, 2022.

\bibitem{beniaguev2021single}
David Beniaguev, Idan Segev, and Michael London.
\newblock Single cortical neurons as deep artificial neural networks.
\newblock {\em Neuron}, 109(17):2727--2739, 2021.

\bibitem{kay2024ant}
Tomas Kay, Alba Motes-Rodrigo, Arthur Royston, Thomas~O Richardson, Nathalie Stroeymeyt, and Laurent Keller.
\newblock Ant social network structure is highly conserved across species.
\newblock {\em Proceedings B}, 291(2027):20240898, 2024.

\bibitem{richardson2021leadership}
Thomas~O Richardson, Andrea Coti, Nathalie Stroeymeyt, and Laurent Keller.
\newblock Leadership--not followership--determines performance in ant teams.
\newblock {\em Communications biology}, 4(1):535, 2021.

\bibitem{stockmaier2021infectious}
Sebastian Stockmaier, Nathalie Stroeymeyt, Eric~C Shattuck, Dana~M Hawley, Lauren~Ancel Meyers, and Daniel~I Bolnick.
\newblock Infectious diseases and social distancing in nature.
\newblock {\em Science}, 371(6533):eabc8881, 2021.

\bibitem{leckie2024architectural}
Luke Leckie, Mischa~Sinha Andon, Katherine Bruce, and Nathalie Stroeymeyt.
\newblock Architectural immunity: ants alter their nest networks to prevent epidemics.
\newblock {\em bioRxiv}, pages 2024--08, 2024.

\bibitem{sole2019liquid}
Ricard Sol{\'e}, Melanie Moses, and Stephanie Forrest.
\newblock Liquid brains, solid brains, 2019.

\bibitem{muller2020power}
Caroline M{\"u}ller, Barbara~A Caspers, J{\"u}rgen Gadau, and Sylvia Kaiser.
\newblock The power of infochemicals in mediating individualized niches.
\newblock {\em Trends in Ecology \& Evolution}, 35(11):981--989, 2020.

\bibitem{clark2020niche}
Andrew~D Clark, Dominik Deffner, Kevin Laland, John Odling-Smee, and John Endler.
\newblock Niche construction affects the variability and strength of natural selection.
\newblock {\em The American Naturalist}, 195(1):16--30, 2020.

\bibitem{dennett2009darwin}
Daniel Dennett.
\newblock Darwin's “strange inversion of reasoning”.
\newblock {\em Proceedings of the National Academy of Sciences}, 106(supplement\_1):10061--10065, 2009.

\bibitem{beverley1867darwinian}
Robert~Mackenzie Beverley.
\newblock {\em The Darwinian theory of the transmutation of species}.
\newblock J. Nisbet, 1867.

\bibitem{watson2023collective}
Richard Watson and Michael Levin.
\newblock The collective intelligence of evolution and development.
\newblock {\em Collective Intelligence}, 2(2):26339137231168355, 2023.

\bibitem{frith2010social}
Uta Frith and Chris Frith.
\newblock The social brain: allowing humans to boldly go where no other species has been.
\newblock {\em Philosophical Transactions of the Royal Society B: Biological Sciences}, 365(1537):165--176, 2010.

\bibitem{farrar2002early}
M~Jeffrey Farrar and Lisa Maag.
\newblock Early language development and the emergence of a theory of mind.
\newblock {\em First language}, 22(2):197--213, 2002.

\bibitem{de2021role}
Jill~G de~Villiers.
\newblock The role (s) of language in theory of mind.
\newblock In {\em The neural basis of mentalizing}, pages 423--448. Springer, 2021.

\bibitem{milligan2007language}
Karen Milligan, Janet~Wilde Astington, and Lisa~Ain Dack.
\newblock Language and theory of mind: Meta-analysis of the relation between language ability and false-belief understanding.
\newblock {\em Child development}, 78(2):622--646, 2007.

\bibitem{lombard2023causal}
Marlize Lombard and Peter G{\"a}rdenfors.
\newblock Causal cognition and theory of mind in evolutionary cognitive archaeology.
\newblock {\em Biological Theory}, 18(4):234--252, 2023.

\bibitem{momennejad2022collective}
Ida Momennejad.
\newblock Collective minds: social network topology shapes collective cognition.
\newblock {\em Philosophical Transactions of the Royal Society B}, 377(1843):20200315, 2022.

\bibitem{lau2018discovering}
Tatiana Lau, Hillard~T Pouncy, Samuel~J Gershman, and Mina Cikara.
\newblock Discovering social groups via latent structure learning.
\newblock {\em Journal of Experimental Psychology: General}, 147(12):1881, 2018.

\bibitem{woolley2010evidence}
Anita~Williams Woolley, Christopher~F Chabris, Alex Pentland, Nada Hashmi, and Thomas~W Malone.
\newblock Evidence for a collective intelligence factor in the performance of human groups.
\newblock {\em science}, 330(6004):686--688, 2010.

\bibitem{jara2019theory}
Julian Jara-Ettinger.
\newblock Theory of mind as inverse reinforcement learning.
\newblock {\em Current Opinion in Behavioral Sciences}, 29:105--110, 2019.

\bibitem{ruiz2023inverse}
Jaime Ruiz-Serra and Michael~S Harr{\'e}.
\newblock Inverse reinforcement learning as the algorithmic basis for theory of mind: current methods and open problems.
\newblock {\em Algorithms}, 16(2):68, 2023.

\bibitem{strachan2024testing}
James~WA Strachan, Dalila Albergo, Giulia Borghini, Oriana Pansardi, Eugenio Scaliti, Saurabh Gupta, Krati Saxena, Alessandro Rufo, Stefano Panzeri, Guido Manzi, et~al.
\newblock Testing theory of mind in large language models and humans.
\newblock {\em Nature Human Behaviour}, pages 1--11, 2024.

\bibitem{kiciman2023causal}
Emre K{\i}c{\i}man, Robert Ness, Amit Sharma, and Chenhao Tan.
\newblock Causal reasoning and large language models: Opening a new frontier for causality.
\newblock {\em arXiv preprint arXiv:2305.00050}, 2023.

\bibitem{haidle2010working}
Miriam~No{\"e}l Haidle.
\newblock Working-memory capacity and the evolution of modern cognitive potential: implications from animal and early human tool use.
\newblock {\em Current anthropology}, 51(S1):S149--S166, 2010.

\bibitem{watson2011global}
Richard~A Watson, Rob Mills, and Christopher~L Buckley.
\newblock Global adaptation in networks of selfish components: Emergent associative memory at the system scale.
\newblock {\em Artificial Life}, 17(3):147--166, 2011.

\bibitem{shevlin2019apply}
Henry Shevlin and Marta Halina.
\newblock Apply rich psychological terms in ai with care.
\newblock {\em Nature Machine Intelligence}, 1(4):165--167, 2019.

\end{thebibliography}

\documentclass[8pt]{article}

\usepackage[english]{babel}
\usepackage{geometry}
\usepackage{amsmath}
\usepackage{graphicx}
\usepackage{soul}
\usepackage[colorlinks=true, allcolors=blue]{hyperref}

\title{Artificial Theory of Mind and Self-Guided Social Organisation}
\author{Michael S. Harr\'e, Jaime Ruiz-Serra, Catherine Drysdale}

\begin{document}
\date{}
\maketitle

%\begin{abstract}
%Your abstract.
%\end{abstract}

%\section{Introduction}

One of the challenges artificial intelligence (AI) faces is how a collection of agents coordinate their behaviour to achieve goals that are not reachable by any single agent~\cite{wolpert1999introduction,levin2023bioelectric}. In a recent article by Ozmen et al~\cite{ozmen2023six} this was framed as one of six grand challenges: That AI needs to respect human cognitive processes at the human-AI interaction frontier. We suggest that this extends to the AI-AI frontier and that it should also {\it reflect} human psychology, as it is the only successful framework we have from which to build out. In this extended abstract we first make the case for collective intelligence in a general setting, drawing on recent work from single neuron complexity in neural networks and ant network adaptability in ant colonies. From there we introduce how species relate to one another in an ecological network via niche {\it selection}, niche {\it choice}, and niche {\it conformity} with the aim of forming an analogy with human social network development as new agents join together and coordinate. From there we show how our social structures are influenced by our neuro-physiology, our psychology, and our language. This emphasises how individual people within a social network influence the structure and performance of that network in  complex tasks, and that cognitive faculties such as Theory of Mind play a central role. We finish by discussing the current state of the art in AI and where there is potential for further development of a {\it socially embodied collective artificial intelligence} that is capable of guiding its own social structures.

Biological agents are composed of sub-units such as organs, cells, and molecular networks~\cite{levin2019computational}, in which, for example, individual neurons display computational intelligence, or ``competencies", that can scale freely~\cite{levin2022technological}. Beniaguev et al.\cite{beniaguev2021single} demonstrated that a single cortical neuron requires five to eight layers in a deep neural network to approximate its input-output mapping, and so we may conceptualise a neuron as an artificial agent with complex computational abilities that adapts its connections to other agents. Similarly, ant colonies have conserved social structures and division of labor across species separated by over 100 million years of evolution~\cite{kay2024ant}, with leadership roles in individual ants enhancing collective performance~\cite{richardson2021leadership}. In response to pathogenic threats, ants adapt by reorganizing their social behaviors~\cite{stockmaier2021infectious} and modifying their nests to slow disease spread~\cite{leckie2024architectural}. While neural networks exhibit ``solid brain" connectivity (agent-to-agent connectivity is rigid in time) and ant colonies demonstrate ``liquid brain" dynamics~\cite{sole2019liquid} (agent-to-agent interactions are fluid in time), both systems continuously integrate information and adjust their connectivity for the collective benefit.

There are however differences in how freely new agents connect in these two model systems. The adult brain has limited capacity for neurogenesis so that recruitment, placement, and integration of new cells plays a relatively small role in adult mammals. But in liquid brains new agents can freely join collectives and contribute to the collective outcomes of joint actions. The process by which this occurs in ecological network is detailed in~\cite{muller2020power}, where the interaction between a single plant, for example, and its ecological environment results in individualised niches via three main processes: niche choice, niche conformance, and niche construction. Niche choice occurs when an individual selects environmental conditions that align with its phenotype, while niche conformance involves adjusting its phenotype to suit the environment. Niche construction is the modification of the environment to meet individual needs, which may also impact other species~\cite{clark2020niche}. These processes affect both individual, improving the match between the individual and its ecological network, thereby improving their overall fitness. Effective niche realisation relies on communication between encoders-senders and receivers-decoders using so-called {\it info-chemicals} as the signalling mechanism. In sum, interactions between individual plants and animals, communicated over information carrying channels, shape network relations at all levels in ecological networks.

This is also the case in the context of human social networks, where, as we argue later, Theory of Mind (ToM) has a key role to play. In his review of Darwin's {\it strange inversion of reasoning}, Dennet~\cite{dennett2009darwin} opens with the following: 
\begin{quote}
    Darwin's theory of evolution by natural selection unifies the world of physics with the world of meaning and purpose by proposing a deeply counterintuitive ``inversion of reasoning'' (according to a 19th century critic): ``to make a perfect and beautiful machine, it is not requisite to know how to make it''~\cite{beverley1867darwinian}.
\end{quote}
Here we suggest that this was inverted again by the evolution of our very human capacity to be {\it consciously} aware of ourselves and our social and cultural constructs, and to {\it purposefully} manipulate them to our collective benefit. Take this observation about manipulating inter-cellular communication by Watson and Levin~\cite{watson2023collective}: 
\begin{quote}
    This framework [of collective cellular intelligence] makes a strong prediction: if intercellular signalling (not genes) is the cognitive medium of a morphogenetic individual, it should be possible to exploit the tools of behavioural and neuro-science and learn to read, interpret and re-write its information content in a way that allows predictive control over its behaviour (in this case, growth and form) without genetic changes. 
\end{quote}
A counter question then arises: How do humans {\it read, interpret, and re-write our inter-personal information content} in order to exert predictive control over a social group? Similar to how a scientist \textit{external} to a cell collective manipulates inter-cellular signaling to control collective outcomes, individuals \textit{internal} to a human collective can influence inter-personal behaviors to (re)shape social structures in order to alter group outcomes. In both cases, an agent with a goal-directed psychology manipulates inter-agent relationships---whether inter-cellular or inter-personal---to control higher-level outcomes, be it at the organism or societal scale. And so, as new people join a social group and just as new species form niches, rather than integrating a new person based on mechanisms such as random connectivity or rich-get-richer processes, a dynamic integration occurs whereby the networked agents and the prospective agent selectively adjust relationships to accommodate (or reject) the newcomer. This requires complex psychological abilities in order to do it at the time scale of the phenotype, rather than at the evolutionary time-scale, which we address later in this work. It is achieved via a highly evolved set of  cognitive tools that humans achieve this, as we discuss next.

ToM is the ability to mentally represent the internal states of another person~\cite{frith2010social} and a surprising connection exists between the development of ToM and language. The {\it representational} view suggests that both children~\cite{farrar2002early} and adults use specific grammatical structures to represent complex events and reason from them~\cite{de2021role}. These {\it complement structures} are an example of language-as-cognitive tool that has special utility in representing the mental states of other people, for example they allow for the expression of mistakes, lies, or false beliefs. Complement structures are strong predictors of children's false belief understanding (a canonical test of ToM) in longitudinal studies, and training in complement syntax has been shown to improve children's false belief reasoning. It is interesting to note that in developmental learning there is a stronger effect from language to ToM ability than vice versa~\cite{milligan2007language}.

Language and ToM also strongly interact with our causal reasoning in social settings. Lombard and Gärdenfors~\cite{lombard2023causal} put forward three hypotheses: (1) ToM is a crucial component of causal cognition; (2) The more sophisticated causal cognition becomes, the more it relies on ToM; and (3) The evolution of causal cognition increasingly depends on mental representations of hidden variables. The aspect we wish to focus on here is the causal modelling. At the individual level, recent work by Momennejad~\cite{momennejad2022collective} reviewed neuro-imaging evidence for people neurologically encoding their social network topologies and the sharing of these encodings across a social group. This was also demonstrated in the work of Lau et al.~\cite{lau2018discovering} showing that people are able to integrate information about how agents relate to one another in addition to how they relate to oneself in order to infer social group structures. This allows agents to have a collective causal understanding of how individuals socially interact with one another. Finally, there is also evidence for improved collective intelligence when individuals with higher competencies in ToM are present in a social setting~\cite{woolley2010evidence}. From this we conclude that language, a shared collective understanding of social causal relationships, and our ToM are part of a highly integrated cognitive toolbox that we have developed to understand how we collectively fit together to improve our coordination towards collective goals. We emphasise that the collective goals are also psychological constructs of individuals, and in pursuit of these goals people take advantage of their ability to read, interpret, and even re-write the social connections between themselves and other agents in order to achieve a greater collective outcome than any single agent could achieve by themselves.

Where does artificial intelligence sit with respect to these cognitive tools? Inverse reinforcement learning (IRL) has been suggested as a model for ToM in AI~\cite{jara2019theory,ruiz2023inverse} where an agent infers the preferences, e.g. rewards as a function of state, of another agent based solely on its behaviour. Jara-Ettinger~\cite{jara2019theory} discusses some key limitations of IRL but misses the social network within which the agents are situated and the causal cognitive model of social interactions. Large Language Models (LLMs) have also been a very popular test-bed for ToM in AI~\cite{strachan2024testing}, however it has been shown that while LLMs can replicate causal reasoning in some scenarios, they also have highly unpredictable failure modes~\cite{kiciman2023causal} and there are very few studies on {\it causal social cognition}. There is also no evidence to date that the complete suite of cognitive skills mentioned above have been demonstrated by an LLM. In particular, there are no LLMs that are socially embodied in a communication network whereby an individual agent articulates a collective goal, manipulates the social connections between other agents in order to optimise the network structure in order to achieve that goal, and to do so based on a shared representation of each agent's causal model as even early humans could do~\cite{haidle2010working}. However, there has been some promising advances made in formally progressing the state of the art, for example in~\cite{watson2011global}, where Watson et al. show that
\begin{quote}
    ... when self-interested agents can modify how they are affected by other agents (e.g., when they can influence which other agents they interact with), then, in adapting these inter-agent relationships to maximize their own utility, they will necessarily alter them in a manner homologous with Hebbian learning. Multi-agent systems with adaptable relationships will thereby exhibit the same system-level behaviors as neural networks under Hebbian learning.
\end{quote}
There are many other avenues that are being explored in the artificial context that we cannot cover in this extended abstract, but in our talk at GSO-2025 we will cover some of the major themes that have been contributing to this research and where their strengths and weaknesses lie. In broad strokes our presentation will take a leaf from Shevlin and Halina and urge caution in applying rich psychological terms in AI~\cite{shevlin2019apply}.
%This is a cognitive technology that we argue is likely necessary but not sufficient for people to learn to interpret, share and re-write interpersonal communications. 

%Human cognitive structures,
%Human planning and foresight,
%We cognitively encode our social networks,
%Theory of Mind: Its recent recasting in artificial intelligence and human psychology,
%Language and Theory of Mind: Co-constructors of social meaning

\bibliographystyle{unsrt}
\bibliography{sample}

\end{document}